OpenStudy (anonymous):

@Thomaster @Gorv @Ganeshie8 Which is the first step to solve this equation: 10m = 30 A. Divide 10 by 30 B. Multiply both sides by 10 C. Divide both sides by 10 D. Divide both sides by 30

OpenStudy (gorv):

we need to find m

OpenStudy (gorv):

so we need to keep m alone

OpenStudy (gorv):

but 10 is in multiplication so how can u remove it ??

OpenStudy (anonymous):

Times it by both sides?

OpenStudy (gorv):

no divide it both side by 10

OpenStudy (gorv):

if we divide by 10 thn 10 will cancel 10 right??

OpenStudy (anonymous):

Yea

OpenStudy (gorv):

so what is our first step thn ??

OpenStudy (anonymous):

I think 10 / 10 = 1

OpenStudy (gorv):

yeah so first step is ??

OpenStudy (gorv):

what we did in first step??

OpenStudy (anonymous):

divide 10 by 30?

OpenStudy (anonymous):

i meant 30/10

OpenStudy (gorv):

we divide both side by 10

OpenStudy (gorv):

\[\frac{ 10*m }{ 10 }=\frac{ 30 }{ 10 }\]

OpenStudy (anonymous):

-5y = 75

OpenStudy (gorv):

what we will divide by ?? to make y alone ??

OpenStudy (anonymous):

u divide by 5 ?

OpenStudy (anonymous):

-5

OpenStudy (gorv):

oh wow u r good buddy :)

OpenStudy (anonymous):

Lol thanks

OpenStudy (anonymous):

y – 9 = 50

OpenStudy (gorv):

welcome:)

OpenStudy (anonymous):

Np